A pharmacist adds 10 ml of a 20% w/v solution of a drug to a 500 ml D5W for parenteral infusion. What is the % strength of the drug in the infusion solution?
Answer #1

In this case, let x1= 20 and x2= 0, since the active ingredient dextrose in D5W is not considered in this case.

The % strength of the drug in the infusion solution can be calculated as

xpph (w/v)= x1/500ml D5W) + x2/10ml

=0/500ml D5W) + 20%/ 10ml

( To convert 20% into the quantity of active ingredients , it is multiplied by 10 which is the quantity added).

= 0/500mlD5W + 20/100 * 10/10ml

= 0/500ml D5W) + 2gm/10ml

By direct addition, we get,

=  2gm/510ml D5W)

= 2gm/5.10 * 100ml D5W)

= 0.39g/100ml D5W)

= 0.39% (W/V) % (W/V)

The % strength of drug in infusion solution is 0.39% (W/V).